# Reconstructing Physics-Informed Machine Learning for Traffic Flow Modeling: a Multi-Gradient Descent and Pareto Learning Approach

### v1

@@ 4.2 PIML model based on the microscopic traffic flow model
where $\alpha$: the weight of the loss function that balances the contributions made by the data discrepancy and the physics discrepancy; $\hat{\lambda}$: parameters of the physics that are calibrated beforehand; $\theta$: PUNN’s parameters; $f_{\theta}(\cdot)$: PUNN parameterized by $\theta$; $f_{\hat{\lambda}}(\cdot)$: physics parameterized by $\hat{\lambda}$, which is calibrated prior to training; $N_{O}$: the number of observed data; $N_{C}$: the number of collocation states; $\hat{\mathbf{s}}^{(i)}$: the $i$th observed feature; $\mathbf{s}^{(j)}$: the $j$th collocation feature; $\hat{a}^{(i)}$: the $i$th observed acceleration; $a^{(i)}$: the $i$th acceleration predicted by PUNN; $a^{(j)}$: the $j$th acceleration predicted by PUNN; $a_{phy}^{(j)}$: the $j$th acceleration predicted by the physics. The collocation states are functionally equivalent to auxiliary points, which is used to construct the physics residuals. Readers could refer to Section 4.1 or Section 3.2.2 in Mo et al. [2021].
@@
where $x$ and $v$ denote the positions and velocities of the follower, $T_{i}$ is the time horizon of the $i$th trajectory, and $N_{T}$ is the number of trajectories. Since only the IDM-based physics-informed car-following model is present in the official reproduction code, we denote the IDM-based physics-informed car-following model as the original physics-informed car-following model (OPICF) in our paper. As shown in Table 5, all training methods based on multi-gradient descent perform better than the baseline model, and those using DCGD surpass the model based on TMGD. The relative improvement of training methods based on Multi-Dimensional Genetic Algorithms (MDGAs) compared to the baseline is presented in Table 6. All training methods utilizing DCGD algorithms demonstrate a relative improvement of over 20% in the RMSE of $\hat{x}$ compared to $x$, and a relative improvement of over 20% in the RMSE of $\hat{u}$ compared to $u$. We also provide a series of comparisons between predicted and actual positions for sample trajectories based on various models in Figures 5 to 9.
Table 5: Comparison of MGDAs and linear scalarization for physics-informed car-following model Mo et al. [2021]
@@
Table 6: Relative improvement of training based on MGDAs compared to the baseline
Model Relative improvement of $RMSE(\hat{x},x)$ Relative improvement of $RMSE(\hat{u},u)$
TMGD-PICF 5.5667% 8.0949%
DCGD-CENTER-PICF 28.9277% 24.5173%
DCGD-PROJ-PICF 28.2105% 24.6569%
DCGD-AVER-PICF 27.9524% 24.6453%
@@ 5 Conclusion and discussion
In the first example, the MGDAs-based training approach can achieve a comparable performance compared to the baseline model in most cases. In the second example, all MGDAs-based training approaches bring a considerable performance improvement. From these case studies, we can conclude that using MGDAs to train the physics-informed machine learning model is preferable, as long as they do not lead to a significant performance degradation. This training approach is beneficial as long as they do not negatively impact performance because it eliminates the need for hyperparameter searching.
Possible further directions could be to focus on three areas: (1) How to accelerate multi-gradient descent algorithms. Compared to directly using a stochastic gradient descent algorithm like Adam (Kingma and Ba [2014]) to minimize the linear scalarization loss. The training process based on the reformulated multi-objective optimization will lead to a longer training time per iteration. Especially for TMGD algorithms, in each iteration, we need to find the minimum norm point by solving a convex quadratic programming problem. (2) How to handle hard constraints in the training process. All MGDAs used in this paper are only compatible with unconstrained multi-objective optimization. What if we have to consider other constraints in PIML? Is there another way rather than modeling constraints in the loss function directly as a soft constraint term? A possible way is to extend the framework proposed in Lei et al. [2025a], which presents a framework based the multi-gradient descent that can handle multi-objective optimization problems with convex constraints. (3) Research on multi-task learning suggests that TGMD (Sekitani and Yamamoto [1993]) may not consistently produce a well-balanced Pareto front. Instead, TGMD tends to generate a discrete and sparse representation of the Pareto front. Therefore, further studies could explore methods to create a more complete or continuous Pareto front by following the approaches outlined by Lin et al. [2019] and Ma et al. [2020].
@@ 6.1 Proof theorem 5
Suppose $\theta^{*}$ is a minimizer of the scalarized total loss for some $\beta\in[0,1]$. Then, at this point, we have:
$\nabla\mathcal{L}_{\beta}(\theta^{*})=\beta\nabla\mathcal{L}_{1}(\theta^{*})+(1-\beta)\nabla\mathcal{L}_{2}(\theta^{*})=0.$ (34)
This implies that $\nabla\mathcal{L}_{1}(\theta^{*})$ and $\nabla\mathcal{L}_{2}(\theta^{*})$ are linearly dependent and in opposite directions, or one of them is equal to 00. Thus, there exists no common descent direction that strictly reduces both losses simultaneously, implying that $\theta^{*}$ is Pareto stationary. By construction, MGDA identifies Pareto stationary points by minimizing the norm of convex combinations of gradients. Hence, MGDA can also identify and achieve such points. Therefore, every scalarized minimizer can theoretically be reached by MGDA.
@@
Consider a Pareto stationary point $\hat{\theta}$ found by MGDA, satisfying:
$\min_{\alpha\in[0,1]}\|\alpha\nabla\mathcal{L}_{1}(\hat{\theta})+(1-\alpha)\nabla\mathcal{L}_{2}(\hat{\theta})\|^{2}=0.$ (35)
Suppose, for example, that $\nabla\mathcal{L}_{1}(\hat{\theta})=0$ and $\nabla\mathcal{L}_{2}(\hat{\theta})\neq 0$. Then, for any $\beta\in[0,1)$,
$\beta\nabla\mathcal{L}_{1}(\hat{\theta})+(1-\beta)\nabla\mathcal{L}_{2}(\hat{\theta})=(1-\beta)\nabla\mathcal{L}_{2}(\hat{\theta})\neq 0,$ (36)
and only for $\beta=1$ does the combination vanish. However, at $\beta=1$, the scalarized loss reduces to $\mathcal{L}_{1}$ alone, and $\hat{\theta}$ is not necessarily a minimizer of $\mathcal{L}_{2}$. Thus, $\hat{\theta}$ cannot be obtained as a minimizer of any scalarized loss with $\beta\in[0,1]$, except possibly at the endpoints, and in general is not a minimizer for both losses. Therefore, there exist Pareto stationary points found by MGDA that cannot be obtained by optimizing any linear scalarization.
@@ 7 Acknowledgement
This research is supported by the award ”CAREER: Physics Regularized Machine Learning Theory: Modeling Stochastic Traffic Flow Patterns for Smart Mobility Systems (# 2234289)” which is funded by the National Science Foundation.

### v3

@@ 4.2 PIML model based on the microscopic traffic flow model
[…]
@@
where $x$ and $v$ denote the positions and velocities of the follower, $T_{i}$ is the time horizon of the $i$th trajectory, and $N_{T}$ is the number of trajectories. Since only the IDM-based physics-informed car-following model is present in the official reproduction code, we denote the IDM-based physics-informed car-following model as the original physics-informed car-following model (OPICF) in our paper. It is noted that we adopted the same early stop mechanism as Mo et al. [2021] to prevent overfitting, and have the same maximum training iterations. As shown in Table 6, we conduct a series of comparative experiments involving the benchmark training setup, traditional multi-gradient descent Sener and Koltun [2018], dual cone gradient descent Hwang and Lim [2024], dynamic weight averaging (DWA) Liu et al. [2019], and uncertainty weighting (UW) Kendall et al. [2018]. The key distinction among these experiments lies in the choice of loss formulation and training strategy. Specifically, the benchmark, DWA, and uncertainty weighting approaches fall under single-objective optimization frameworks, as they scalarize multiple loss terms into a single objective. In contrast, multi-gradient descent and dual cone gradient descent are genuine multi-objective optimization methods that aim to balance task-specific losses via principled directional updates grounded in multi-task optimization theory. Table 5 summarizes the loss functions and update directions used in a subset of the experimental settings. In this table, $\mathcal{L}_{1}$ denotes the data loss and $\mathcal{L}_{2}$ denotes the physics loss, $k$ is the iteration index, $\sigma_{t}$ denotes the noise associated with the $t$th loss term, and $\alpha^{*}$ is the a weight vector obtained from Problem 12. While the table does not explicitly list the formulation for dual cone gradient descent (DCGD), it similarly treats $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ as two distinct objectives within a multi-objective framework. Its update direction is computed via a different strategy, as detailed in Section 3.1.
Table 5: Comparison of different training set up
Model Loss function Update direction
OPICF $\mathcal{L}=0.9\cdot\mathcal{L}_{1}+0.1\cdot\mathcal{L}_{2}$ $\mathbf{d}=-0.9\cdot\nabla\mathcal{L}_{1}-0.1\nabla\mathcal{L}_{2}$
DWA-PICF $\mathcal{L}=\sum_{t=1}^{2}w_{t}^{(k)}\mathcal{L}_{t},w_{t}^{(k)}=\frac{2\exp{\bigg(\frac{\mathcal{L}_{t}^{(k-1)}}{\mathcal{L}_{t}^{(k-2)}}/2\bigg)}}{\sum_{j=1}^{2}\exp{\bigg(\frac{\mathcal{L}_{j}^{(k-1)}}{\mathcal{L}_{j}^{(k-2)}}/2\bigg)}}$ $\mathbf{d}=-\sum_{t=1}^{2}w_{t}^{k}\nabla\mathcal{L}_{t}$
UW-PICF $\mathcal{L}=\sum_{t=1}^{2}\bigg(\frac{1}{2\sigma_{t}^{2}}\mathcal{L}_{t}\bigg)+\log\sigma_{t}$ $\mathbf{d}=-\sum_{t=1}^{2}\frac{1}{2\sigma_{t}^{2}}\nabla\mathcal{L}_{t}$
TMGD-PICF $\mathcal{L}_{1},\mathcal{L}_{2}$ $\mathbf{d}=-\sum_{t=1}^{2}\alpha^{*}_{t}\nabla\mathcal{L}_{t}$
All training methods based on multi-gradient descent outperform the baseline model as well as the DWA and UW strategies. Among them, the DCGD variants consistently yield the best results. Specifically, models trained using DCGD-CENTER, DCGD-PROJ, and DCGD-AVER achieve significantly lower RMSEs in both position and speed compared to scalarization-based methods and earlier MGDA baselines.
As shown in Table 6, the DCGD-CENTER-PICF model achieves the lowest position error, with an RMSE of 5.7275 $\pm$ 0.2051, while DCGD-PROJ-PICF achieves the best velocity accuracy, with an RMSE of 0.6478 $\pm$ 0.0997. The performance differences among the three DCGD variants are relatively minor and not statistically significant based on standard deviation overlaps. However, the improvement over scalarization-based models such as OPICF (with RMSE of 8.0587 $\pm$ 4.0024 in position and 0.8598 $\pm$ 0.1357 in speed) is substantial and consistent, indicating the effectiveness of direction correction in multi-objective optimization.
We also provide visual comparisons between predicted and actual trajectories for each training method in Figures 7 to 11. Each method was evaluated on 10 independent test trajectories. In these figures, the red dashed line represents the ground-truth trajectory, while the blue line denotes the predicted trajectory generated by the physics-informed car-following model trained with each method. As illustrated, DCGD-based models produce predictions that more closely track the true trajectory, both in position and dynamic pattern, further supporting the quantitative findings.
Table 6: Performance comparison between linear scalarization-based and MGDA-based training for physics-informed car-following models
@@
Figure 8: Comparison of predicted and real positions for ten Sample trajectories based on UW-PICF
Figure 9: Comparison of predicted and real positions for ten Sample trajectories based on DWA-PICF
@@ 5 Conclusion and discussion
In the first example, the MGDAs-based training approach can achieve a comparable performance compared to the baseline model in most cases. In the second example, all MGDAs-based training approaches bring a considerable performance improvement. From these case studies, we can conclude that using MGDAs to train the physics-informed machine learning model is preferable and beneficial over the traditional training method because it eliminates the need for hyperparameter searching that can be extremely tricky and time-consuming. Both traditional multi-gradient descent and dual cone gradient descent methods can theoretically ensure that the trained parameters reach a Pareto stationary point. This suggests that neither method has a theoretical advantage over the other, implying that each method may outperform the other in different situations. We suggest keeping both training methods based on multi-gradient descent to replace the traditional training logic.
Possible further directions could be to focus on three areas: (1) How to accelerate multi-gradient descent algorithms. Compared to directly using a stochastic gradient descent algorithm like Adam (Kingma and Ba [2014]) to minimize the linear scalarization loss. The training process based on the reformulated multi-objective optimization will lead to a longer training time per iteration. Especially for TMGD algorithms, in each iteration, we need to find the minimum norm point by solving a convex quadratic programming problem. (2) How to handle hard constraints in the training process. All MGDAs used in this paper are only compatible with unconstrained multi-objective optimization. What if we have to consider other constraints in PIML? Is there another way rather than modeling constraints in the loss function directly as a soft constraint term? A possible way is to extend the framework proposed in Lei et al. [2025a], which presents a framework based the multi-gradient descent that can handle multi-objective optimization problems with convex constraints. (3) Research on multi-task learning suggests that TGMD (Lin et al. [2019]) may not consistently produce a well-balanced Pareto front. Instead, TGMD tends to generate a discrete and sparse representation of the Pareto front. Therefore, further studies could explore methods to create a more complete or continuous Pareto front by following the approaches outlined by Lin et al. [2019] and Ma et al. [2020].
@@ 6.1 Proof theorem 5
Suppose $\theta^{*}$ is a (local) minimizer of the scalarized total loss for some $\beta\in(0,1)$. Then
$\nabla\mathcal{L}_{\beta}(\theta^{*})\;=\;\beta\nabla\mathcal{L}_{1}(\theta^{*})+(1-\beta)\nabla\mathcal{L}_{2}(\theta^{*})\;=\;0$ (34)
Equivalently, $0\in\mathrm{co}\{\nabla\mathcal{L}_{1}(\theta^{*}),\nabla\mathcal{L}_{2}(\theta^{*})\}$, hence $\theta^{*}$ is Pareto stationary. By Theorem 3 and 4, MGDA identifies such points as fixed points without common strict descent. Therefore, every interior scalarized minimizer can be reached (identified) by MGDA.
@@
Consider a Pareto stationary point $\hat{\theta}$ (for example, one found by MGDA) with
$\nabla\mathcal{L}_{1}(\hat{\theta})=0\qquad\nabla\mathcal{L}_{2}(\hat{\theta})\neq 0$ (35)
For any $\beta\in(0,1)$ we have
$\nabla\mathcal{L}_{\beta}(\hat{\theta})=\beta\nabla\mathcal{L}_{1}(\hat{\theta})+(1-\beta)\nabla\mathcal{L}_{2}(\hat{\theta})=(1-\beta)\nabla\mathcal{L}_{2}(\hat{\theta})\neq 0$ (36)
Thus $\hat{\theta}$ is not a stationary point, and therefore it is not a minimizer of any interior linear scalarization. Consequently, certain Pareto stationary points found by MGDA cannot be obtained by optimizing any interior linear scalarization. ∎
@@ 7 Acknowledgement
This research is supported by the award ”CAREER: Physics Regularized Machine Learning Theory: Modeling Stochastic Traffic Flow Patterns for Smart Mobility Systems (# 2234289)” which is funded by the National Science Foundation. We would like to thank Professor Dong-Young Lim from Ulsan National Institute of Science and Technology for generously sharing the original vector versions of the figures for reproduction.